Tower Corp. had the following stock outstanding and Retained Earnings at December 31, 2018:

   

Common Stock (par $8; outstanding, 33,000 shares)$264,000

Preferred Stock, 8% (par $10; outstanding, 6,300 shares) 63,000

Retained Earnings 283,000

On December 31, 2018, the board of directors is considering the distribution of a cash dividend to the common and preferred stockholders. No dividends were declared during 2016 or 2017, and none have been declared yet in 2018. Three independent cases are assumed:

Case A:The preferred stock is noncumulative; the total amount of 2018 dividends would be $12,900.

Case B:The preferred stock is cumulative; the total amount of 2018 dividends would be $15,120. Dividends were not in arrears prior to 2016.

Case C:Same as Case B, except the total dividends are $69,000.

Required:

1-a. Compute the amount of 2018 dividends, in total that would be payable to each class of stockholders if dividends were declared as described in each case. TIP: Preferred stockholders with cumulative dividends are to be paid dividends for any prior years (in arrears) and for the current year before common stockholders are paid.

1-b. Compute per case, the 2018 dividends per share, payable to each class of stockholders.
Answer #1

Preference Dividend for year = $63000 x 8% = $5040

Case A
Preference Dividend = $5040, Per share = $0.80 per share
Common Dividend = $12900-5040 = $7860. Per share = $0.24 per share

Case B
Preference Dividend = $5040x3 = $15120, Per share = $2.40 per share
Common Dividend = $15120-15120 = $0. Per share = $0 per share

Case C
Preference Dividend = $5040x3 = $15120, Per share = $2.40 per share
Common Dividend = $69000-15120 = $53880. Per share = $1.63 per share